# MCVO: A Generic Visual Odometry for Arbitrarily Arranged Multi-Cameras

### v1

@@ Abstract
Making multi-camera visual SLAM systems easier to set up and more robust to the environment is always one of the focuses of vision robots. Existing monocular and binocular vision SLAM systems have narrow FoV and are fragile in textureless environments with degenerated accuracy and limited robustness. Thus multi-camera SLAM systems are gaining attention because they can provide redundancy for texture degeneration with wide FoV. However, current multi-camera SLAM systems face massive data processing pressure and elaborately designed camera configurations, leading to estimation failures for arbitrarily arranged multi-camera systems. To address these problems, we propose a generic visual odometry for arbitrarily arranged multi-cameras, which can achieve metric-scale state estimation with high flexibility in the cameras’ arrangement. Specifically, we first design a learning-based feature extraction and tracking framework to shift the pressure of CPU processing of multiple video streams. Then we use the rigid constraints between cameras to estimate the metric scale poses for robust SLAM system initialization. Finally, we fuse the features of the multi-cameras in the SLAM back-end to achieve robust pose estimation and online scale optimization. Additionally, multi-camera features help improve the loop detection for pose graph optimization. Experiments on KITTI-360 and MultiCamData datasets validate the robustness of our method over arbitrarily placed cameras. Compared with other stereo and multi-camera visual SLAM systems, our method obtains higher pose estimation accuracy with better generalization ability. Our codes and online demos are available at https://github.com/JunhaoWang615/MCVO.
@@ I Introduction
Visual Simultaneous Localization and Mapping (vSLAM) is a fundamental technique in robotics and autonomous navigation, enabling a system to estimate its motion relative to its environment using visual sensors. Traditional monocular or stereo SLAM systems often require strict camera configurations and rely heavily on integrating inertial measurement units (IMUs) for true scale pose estimation and map construction [1, 2]. However, these approaches are limited by the narrow FoV and perform poorly in environments where camera placement flexibility is paramount, such as most cars with 6 different cameras. Recently, multi-camera visual SLAM allows for greater robustness in these situations, which can cover more surrounding visual scenes to provide redundancy for poorly textured environments. It opens up new possibilities for applications such as UAV navigating through cluttered environments [3], and autonomous driving with surrounding multi-cameras [4].
Figure 1: Illustration of the proposed MCVO system. An example of using two fisheye and one pinhole camera on KITTI-360 dataset [5] for state estimation. The proposed MCVO obtains better performance compared with ORB-SLAM3 [2] using the front stereo camera.
@@
To address these challenges, we propose a generic multi-camera visual odometry system, i.e., MCVO, which takes only multiple rigidly bundled cameras at arbitrary orientations and positions. We first analyze the state-of-the-art feature association methods from the accuracy-efficiency perspective and design a learning-based feature extraction and tracking framework to shift the computation pressure of CPU processing of multiple video streams. Then, we initialize the SLAM system to obtain true scale body poses based on the rigid constraint between the aligned poses for each camera using SfM. In the backend, we fuse multi-camera features to achieve robust pose estimation and scale optimization. The multi-camera features are further concatenated in Bag Of Words (BoW) for loop closure detection. Through rigorous testing and practical implementation on the KITTI-360 and MultiCamData datasets, we aim to demonstrate the efficacy of our system in enhancing the capabilities of autonomous agents across various domains with unprecedented levels of flexibility and generalizability. The highlights of the proposed system: (i) It enables the use of multiple cameras positioned in any orientation with only the requirement of the extrinsic parameters. (ii) By eliminating the dependency on IMU and providing the flexibility to handle cameras in arbitrary configurations, it focuses solely on visual information, no matter overlapping or non-overlapping cameras, which allows for true scale estimation and online optimization, leading to enhanced accuracy and robustness. (iii) Besides, it can accommodate a variety of camera types, including but not limited to fisheye and standard pinhole cameras, making it suitable for a broad range of applications. The main contributions are listed as follows:
  • We propose a generic visual SLAM system framework for arbitrarily arranged multi-cameras, with novel designs for front-end multi-camera feature detection and matching, system initialization, back-end optimization, and loop closure.
  • We comprehensively consider the existing SOTA feature association methods and design a SuperPoint-based extractor and LK tracking front-end, shifting the computational pressure of the CPU and improving the stability of feature matching, which finally provides a guarantee for the stability and efficiency of multi-eye SLAM.
  • We propose a SLAM scale estimation strategy based on the trajectory consistency of multi-cameras, which is compatible with different camera models and with or without camera FoV overlap, demonstrating strong generalization and stability.
@@ II-A Efficient Feature Association Frontend
To enhance system efficiency, some approaches utilize GPUs for parallel processing, shifting the computationally intensive front-end of multi-camera systems to reduce CPU usage and latency. Common feature detectors, such as Shi-Tomasi and Harris, have been implemented in CUDA Visual Library (VILIB) [11] for fast feature extraction. He et al. [9] developed front-end feature tracking algorithms that run on GPUs using NVIDIA’s VPI. Additionally, learning-based methods for feature extraction and tracking [12][13] have been proposed to enhance feature robustness while reducing the computational burden on CPU. Pandey et al. [14] introduced a visual odometry algorithm combining deep learning with optical flow tracking.
@@
II-B Multiple Camera Visual Odomotry
@@ III-A Overview
Our proposed generic multi-camera visual odometry framework is illustrated in Fig. 2. The main inputs to the framework are the synchronized multi-camera video sequences. The multiple cameras are rigidly bundled and calibrated in advance with known intrinsic and extrinsic parameters. The output is the metric scale 6-DoF robot poses in the real-world environment. The pipeline comprises four components: frontend feature extraction, pose and metric scale initialization, backend optimization, and loop closure. To accelerate the frontend of multi-camera feature association, we employ feature extraction with GPU acceleration and 3-priority feature selection (Section III-B. Then we initialize the multi-camera system of metric scale body poses using the multi-camera poses and extrinsic parameters (Section III-C). To ensure the realism of the motion scale, we perform an adaptive correlation of the scale deviations during the backend optimization (Section III-D). Given the larger FoV of the multi-camera system, we design a more robust multi-camera omnidirectional loop detection algorithm (Section III-E). We further optimize the body poses in the loop with the pose graph constraint.
@@ III-B Efficient Multi-Camera Frontend
The frontend of our proposed MCVO primarily processes multi-camera images, extracts and matches useful features, and then provides reliable landmarks for the backend optimization.
Improving feature detection quality and association efficiency are essential for the visual odometry frontend. As the number of cameras increases, the volume of extracted and tracked features grows significantly, leading to a considerable computational burden for real-time performance. Regarding these issues, we utilize SuperPoint [30] for feature extraction and LK optical flow for feature tracking. This process is accelerated via GPU, substantially reducing CPU usage brought by multiple cameras and enhancing performance in handling image streams.
We further conduct a feature selection process, to reduce the CPU load during optimization and improve estimation performance. The traditional quad-tree algorithm [31] tends to place new features close to existing features in texture-rich regions, which might reduce the optimization accuracy due to insufficient constraints. Our feature selection algorithm avoids excessive feature concentration to ensure an even distribution of features through three key priorities.
@@
    Quantity priority: Prioritizes nodes with more features to enhance processing efficiency.
@@
    Tracked feature priority: Prioritizes reducing the density of points in texture-rich areas by avoiding adding useless features near tracked points.
@@
    High score feature priority: Prioritizes features with higher scores.
@@ III-C Multi-Camera Initialization
The initialization of our MCVO system aims to estimate initial poses and determine the metric scale of motions. This process includes the pose initialization for each camera and metric scale estimation for body poses.
We first calculate the parallax for each camera in a 10-frame window. When all the camera frames at time $t$ have sufficient parallax (greater than 30 pixels), we set the world coordinate origin at time $t$ for all cameras. The camera with the most tracking features at time $t$ is selected as the principal camera. Simultaneously, we estimate the poses for each camera independently using Structure from Motion (SfM) on the landmarks provided by the frontend for the 10 frames in the initialization window. It should be noted that the SfM is conducted on 10 frames of each monocular camera, therefore, the scale of each camera is not on the same metric level.
Subsequently, given the poses for all cameras, we first calculate the body frame (principal camera frame) poses using the camera extrinsic parameter
$\bm{p}^{b_{c}}_{t}=\bm{R}^{c}_{t}\bm{p}^{c}+s^{c}\bm{p}^{c}_{t},$ (1)
where $\bm{R}^{c}_{t}$, $\bm{p}^{c}_{t}$ are the estimated rotation and translation for camera $c$ at timestamp $t$, $s^{c}$ is the scale parameter for camera $c$ to be estimated, $\bm{p}^{c}$ is the translational extrinsic parameter of camera $c$ and $\bm{P}^{b_{c}}_{t}$ is the body frame translation estimated from camera $c$ at timestamp $t$.
However, because each camera’s pose is computed independently, their scales are not aligned, resulting in significant discrepancies in body poses. Theoretically, the estimated body pose should be consistent across different cameras. Therefore, we take the differences between body poses calculated by different cameras as the residuals and define the optimization problem as follows:
$\min_{s_{1},s_{2},...,s_{n_{c}}}\sum_{t}\sum_{i=1}^{n_{c}}\sum_{j=i+1}^{n_{c}}$ $||\bm{p}^{b_{i}}_{t}-\bm{p}^{b_{j}}_{t}||^{2},$ (2)
where $n_{c}$ is the number of cameras, $s_{k}(k\in[1,n_{c}])$ is the scale parameter for camera $k$, $\bm{P}^{b_{c}}_{t}$ is the body frame translation estimated from camera $k$ at timestamp $t$. To solve the least squares problem for Eq. 2, we use the Levenberg-Marquardt algorithm from the Ceres Solver library [32].
We readjust camera poses according to the estimated metric scale, then re-triangulate all features using the updated poses to obtain their metric depths. After the initial estimation of poses and scales, we proceed with a sliding window-based backend optimization for state estimation.
@@ III-D Scale Correction and Backend Optimization
After initializing the pose estimation system with metric scale. We use a vision-only bundle adjustment with the extracted features from all cameras to minimizes the sum of prior and the Mahalanobis norm of all visual measurement residuals to obtain a maximum posterior state estimation,
$\min_{\mathcal{X}}$$\Big{\{}\|\bm{r}_{0}\|^{2}_{\mathbf{\Sigma}_{0}}+\sum_{i\in\mathcal{C}}\sum_{j\in\mathcal{F}_{i}}{\bm{\rho}(\|\bm{r}_{v_{(i,j)}}\|}^{2}_{\mathbf{\Sigma}_{{v}_{(i,j)}}})\Big{\}}$ (3)
$\mathcal{X}$ $=[\bm{x}_{1}^{b},\bm{x}_{2}^{b},...,\bm{x}_{t}^{b},\bm{\lambda}_{1},\bm{\lambda}_{2},...,\bm{\lambda}_{n_{c}}]$
where the state vector $\mathcal{X}$ includes the body poses $\bm{x}_{t}^{b}$ in the sliding window and all the feature depths $\bm{\lambda}_{c}$ observed by the $c$-th camera, $\bm{\rho}(\cdot)$ is the robust huber loss, $\mathcal{C}$ is the set of cameras, $\mathcal{F}_{i}$ is the set of reliable features in $i$-th camera, $\bm{r}_{0}$ is the marginalization factor, $\bm{r}_{v_{(i,j)}}$ is the $j$-th landmark reprojection residual in the $i$-th camera. The definition of all the residuals can be found in [1]. The Dogleg algorithm in the Ceres Solver [32] is employed to solve this least squares problem.
Scale Correction. The optimization in Eq. 3 continuously provides accurate body poses and feature depths. However, scale error gradually accumulates as the map expands. To address this, we conduct scale correction alongside the backend optimization at regular intervals. During scale correction, a vision-only bundle adjustment is first performed for each camera to determine its poses independently of the body. We then apply the scale estimation algorithm described in Section III-C to obtain the scale factors $\bm{s}$. Instead of directly adjusting the pose scales, we modify the feature depths in each camera, allowing for a smoother scale correction of poses through subsequent state estimation.

### v2

@@ Abstract
Making multi-camera visual SLAM systems easier to set up and more robust to the environment is attractive for vision robots. Existing monocular and binocular vision SLAM systems have narrow sensing Field-of-View (FoV), resulting in degenerated accuracy and limited robustness in textureless environments. Thus multi-camera SLAM systems are gaining attention because they can provide redundancy with much wider FoV. However, the usual arbitrary placement and orientation of multiple cameras make the pose scale estimation and system updating challenging. To address these problems, we propose a robust visual odometry system for rigidly-bundled arbitrarily-arranged multi-cameras, namely MCVO, which can achieve metric-scale state estimation with high flexibility in the cameras’ arrangement. Specifically, we first design a learning-based feature tracking framework to shift the pressure of CPU processing of multiple video streams to GPU. Then we initialize the odometry system with the metric-scale poses under the rigid constraints between moving cameras. Finally, we fuse the features of the multi-cameras in the back-end to achieve robust pose estimation and online scale optimization. Additionally, multi-camera features help improve the loop detection for pose graph optimization. Experiments on KITTI-360 and MultiCamData datasets validate its robustness over arbitrarily arranged cameras. Compared with other stereo and multi-camera visual SLAM systems, our method obtains higher pose accuracy with better generalization ability. The code will also be made publicly available soon.
@@ I Introduction
Visual Simultaneous Localization and Mapping (vSLAM) is a fundamental technique in robotics and autonomous navigation, enabling a system to estimate its motion relative to its environment using visual sensors. Traditional monocular or stereo SLAM systems often require strict camera configurations and rely heavily on integrating inertial measurement units (IMUs) for metric-scale pose estimation and map construction [1, 2]. However, these approaches are limited by the narrow FoV and perform poorly in environments where camera placement flexibility is paramount, such as most cars with 6 cameras. Recently, multi-camera visual SLAM allows for greater robustness in these situations, which can cover more surrounding visual scenes to provide redundancy for poorly textured environments. It opens up new possibilities for applications such as UAV navigating through cluttered environments [3], and autonomous driving with surrounding multi-cameras [4].
Figure 1: Illustration of the proposed MCVO system. An example of using 3-cam and 4-cam setups on KITTI-360 dataset [5] for state estimation. The proposed MCVO performs better than ORB-SLAM3 [2] using the front stereo camera.
@@
To address these challenges, we propose a robust multi-camera visual odometry system, i.e., MCVO, which takes only multiple rigidly bundled cameras at arbitrary orientations and positions, outputting high-precision metric-scale body poses. We first design a learning-based feature extraction and tracking framework to shift the computation pressure of CPU to GPU. Then, we initialize the SLAM system to obtain true scale body poses based on the rigid constraint between the aligned poses for each camera using SfM. In the backend, we fuse multi-camera features to achieve robust pose estimation and scale optimization. The multi-camera features are further concatenated in Bag Of Words (BoW) for loop closure detection. Through rigorous testing and practical implementation on the KITTI-360 and MultiCamData datasets, we aim to demonstrate the effectiveness of our MCVO system in enhancing the capabilities of autonomous agents across various domains with unprecedented levels of flexibility and generalizability. The highlights of the proposed system: (i) It enables the use of multiple cameras positioned in any orientation with only the requirement of the extrinsic parameters. (ii) By eliminating the dependency on IMU and providing the flexibility to handle cameras in arbitrary configurations, it focuses solely on visual information, no matter overlapping or non-overlapping cameras, which allows for true scale estimation and online optimization, leading to enhanced accuracy and robustness. (iii) Besides, it can accommodate a variety of camera types, including but not limited to fisheye and standard pinhole cameras, making it suitable for a broad range of applications. The main contributions are listed as follows:
  • We propose a generic visual SLAM system framework based on universal multi-camera topology, which innovatively incorporates multi-camera feature extraction frontend, metric-scale pose initialization and optimization, and multi-camera loop closure.
  • We pioneer a robust odometry scale estimation strategy for arbitrarily arranged multi-cameras. Through establishing multi-camera manifold consistency constraints on motion trajectories, we effectively calculate metric-scale poses for multi-camera systems in the initialization. Subsequently, we continuously refine metric scale in the backend optimization.
  • We develop a cascaded tracking algorithm integrating SuperPoint features and Lucas-Kanade optical flow, effectively reducing computational load while enhancing loop closure through multi-camera collaborative verification.
@@ II-A Efficient Feature Association Frontend
To enhance system efficiency, some approaches utilize GPUs for parallel processing, shifting the computationally intensive front-end of multi-camera systems to reduce CPU usage and latency. Common feature detectors, such as Shi-Tomasi and Harris, have been implemented in CUDA Visual Library (VILIB) [11] for fast feature extraction. He et al. [9] developed front-end feature tracking algorithms that run on GPUs using NVIDIA’s VPI. Additionally, learning-based methods for feature extraction and tracking [12][13] have been proposed to enhance feature robustness while reducing the computational burden on the CPU. Pandey et al. [14] introduced a visual odometry algorithm combining deep learning with optical flow tracking.
@@
II-B Multiple Camera Visual Odometry
@@ III-A Overview
Our proposed multi-camera visual odometry framework is illustrated in Fig. 2. The main inputs to the framework are the synchronized multi-camera video sequences. The multiple cameras are rigidly bundled and calibrated in advance with known intrinsic and extrinsic parameters. The output is the metric scale 6-DoF robot poses in the real-world environment. The pipeline comprises four components: frontend feature extraction, pose and metric scale initialization, backend optimization, and loop closure. To accelerate the frontend of multi-camera feature association, we employ feature extraction with GPU acceleration and 3-priority feature selection. Then we initialize the multi-camera system with metric-scale poses using the motion invariance constraints and extrinsic parameters. To ensure the realism of the motion scale, we perform an adaptive correlation of the scale deviations during the backend optimization. Given the larger FoV of the multi-camera system, we design a more robust multi-camera omnidirectional loop detection algorithm. We further optimize the body poses in the loop with the pose graph constraint.
@@ III-B Efficient Multi-Camera Frontend
The frontend of our proposed MCVO primarily processes multi-camera images, extracts and matches useful features, and then provides reliable landmarks for system initialization and backend optimization.
Enhancing feature detection quality and association efficiency is critical for the visual odometry frontend. With an increasing number of cameras, the volume of extracted and tracked features grows exponentially, imposing a substantial computational burden on the edge-computing device. To address these challenges, we implement SuperPoint [30] for feature extraction and LK optical flow for tracking. Leveraging GPU acceleration, this framework significantly reduces CPU overhead from multi-camera setups while improving image stream processing efficiency.
To mitigate CPU load during optimization and enhance estimation accuracy, we propose a feature selection algorithm, the 3-priority quad-tree, which ensures uniform feature distribution. The conventional quadtree method clusters feature in high-texture regions, resulting in insufficient geometric constraints and compromised backend optimization accuracy. In contrast, our approach enforces spatial diversity through three criteria to mitigate feature over-concentration.
@@
    Quantity priority: Prioritize nodes with more features to enhance processing efficiency.
@@
    Tracked feature priority: Prioritize reducing the density of points in texture-rich areas by avoiding adding useless features near tracked points.
@@
    High score feature priority: Prioritize features with higher scores.
@@ III-C Multi-Camera Initialization
The initialization phase of our MCVO system addresses two critical objectives: establishing initial camera poses and recovering the metric scale of motion trajectories. This process involves two components: camera pose initialization and cross-sensor metric scale calibration.
Our approach initiates by analyzing feature parallaxes within a 10-frame sliding window across all cameras. When the parallax of all camera streams exceeds 30 pixels at timestamp $t$, we establish the world coordinate system anchored at this timestamp. The camera exhibiting the highest feature tracking stability at $t$ is designated as the principal camera. Concurrent with this spatial anchoring, we perform independent Structure-from-Motion (SfM) reconstructions for each camera using their respective observed landmarks within the initialization window. The pose estimations for each camera $c$ in the sliding window are mathematically represented as $\mathcal{R}^{c}$ and $\mathcal{T}^{c}$, where $\bm{R}^{c}_{t}\in\text{SO}(3)$ denotes the rotation matrix and $\bm{T}^{c}_{t}\in\mathbb{R}^{3\times 1}$ represents the translation vector at frame $t$. Then we have
$\mathcal{R}^{c}$ =$=\left\{\bm{R}^{c}_{t},t\in[0,10]\right\},$ (1)
$\mathcal{T}^{c}$ =$=\left\{\bm{T}^{c}_{t},t\in[0,10]\right\}.$
It should be emphasized that these monocular SfM instances yield scale-ambiguous poses, necessitating subsequent metric scale unification through multi-camera constraints.
Concretely, building upon the camera pose estimations, we can derive the body frame poses (attached to the principal camera) through extrinsic calibration parameters $\bm{r}^{c}$ and $\bm{t}^{c}$. Notably, the scale ambiguity inherent in monocular SfM causes inter-camera scale misalignment, manifesting as inconsistent body pose estimates across sensor streams. To resolve this scale discrepancy, we introduce scale factors $\bm{s}^{c}\in\mathbb{R}^{+}$ for each camera to enable metric-scale body pose estimation:
$\begin{bmatrix}\bm{R}^{b_{c}}_{t}&\bm{T}^{b_{c}}_{t}\\ \bm{0}&\bm{1}\end{bmatrix}$ $=\begin{bmatrix}\bm{R}^{c}_{t}&s^{c}\bm{T}^{c}_{t}\\ \bm{0}&\bm{1}\end{bmatrix}\begin{bmatrix}(\bm{r}^{c})^{T}&-(\bm{r}^{c})^{T}\bm{t}^{c}\\ \bm{0}&\bm{1}\end{bmatrix}$ (2)
$=\begin{bmatrix}\bm{R}^{c}_{t}\bm{(}{r}^{c})^{T}&-\bm{R}^{c}_{t}(\bm{r}^{c})^{T}\bm{t}^{c}+s^{c}\bm{T}^{c}_{t}\\ \bm{0}&\bm{1}\end{bmatrix}$
Leveraging the theoretical motion consistency constraint that body trajectories should be scale-invariant across camera perspectives, we formulate a residual term for camera $i$ and $j$:
$\bm{\hat{e}}(\bm{T}^{b_{i}}_{t},\bm{T}^{b_{j}}_{t})=$ $\bm{T}^{b_{i}}_{t}-\bm{T}^{b_{j}}_{t}$ (3)
$=$ $-\bm{R}^{i}_{t}(\bm{r}^{i})^{T}\bm{t}^{i}+s^{i}\bm{T}^{i}_{t}+\bm{R}^{j}_{t}(\bm{r}^{j})^{T}\bm{t}^{j}-s^{j}\bm{T}^{j}_{t}$
$=$ $\begin{bmatrix}\bm{T}^{i}_{t}&-\bm{T}^{j}_{t}\end{bmatrix}\begin{bmatrix}s^{i}\\ s^{j}\end{bmatrix}+\bm{\theta},$
where $\theta=-\bm{R}^{i}_{t}(\bm{r}^{i})^{T}\bm{t}^{i}+\bm{R}^{j}_{t}(\bm{r}^{j})^{T}\bm{t}^{j}$. Since camera rotation matrices $\bm{R}^{c}_{t}$ are scale-independent, $\bm{\theta}$ is related to the camera extrinsic parameters $[\bm{r}^{c},\bm{t}^{c}]$ and camera rotation matrices $\bm{R}^{c}_{t}$, thus $\bm{\theta}$ will be also scale-independent. Building upon the pairwise residual formulation in (3), we extend the constraint to $N$-camera systems through block matrix construction. Let the scale vector encapsulate all camera scale factors, the residual takes the form:
$\bm{e}(\bm{T}^{b_{i}}_{t},\bm{T}^{b_{j}}_{t})$ $=\begin{bmatrix}\bm{0}_{\alpha}&\bm{T}^{i}_{t}&\bm{0}_{\beta}&-\bm{T}^{j}_{t}&\bm{0}_{\gamma}\end{bmatrix}\bm{s}+\bm{\theta}$ (4)
$=\bm{F}\bm{s}+\bm{\theta},$
$\bm{s}$ $=\begin{bmatrix}s^{1}&s^{2}&\ldots&s^{N}\end{bmatrix}^{T}$
where $\alpha=3\times(i-1)$, $\beta=3\times(j-i-1)$, $\gamma=3\times(N-j),(i<j)$. $\bm{F}\in\mathbb{R}^{3\times N}$ forms the sparse observation matrix. This construction preserves the problem’s sparsity pattern for efficient large-scale optimization. $\bm{\theta}$ denotes the $\bm{s}$-independent component in Equation (3). The complete scale estimation problem then reduces to solving the constrained least-squares minimization:
$\min_{s_{1},s_{2},...,s_{N}}\sum_{t}\sum_{i=1}^{N}\sum_{j=i+1}^{N}$ $||\bm{e}(\bm{T}^{b_{i}}_{t},\bm{T}^{b_{j}}_{t})||^{2}$ (5)
The scale estimation is formulated as a manifold-aware optimization problem within the ceres solver [31], with scale factor $\bm{s}$ parameterized over Lie groups. We solve the optimization problem using the Levenberg-Marquardt algorithm, which effectively handles the sparse Jacobian structure, with automatic differentiation ensuring numerical precision in partial derivative calculations.
@@ III-D Scale Correction and Backend Optimization
Following metric scale initialization, we formulate a multi-camera visual bundle adjustment problem that jointly optimizes camera poses and 3D landmarks by minimizing the Mahalanobis-weighted reprojection residuals:
$\min_{\mathcal{X}}$$\Big{\{}\|\bm{r}_{0}\|^{2}_{\mathbf{\Sigma}_{0}}+\sum_{c\in N}\sum_{f\in\mathcal{F}_{c}}{\bm{\rho}(\|\bm{r}_{v_{(c,f)}}\|}^{2}_{\mathbf{\Sigma}_{{v}_{(c,f)}}})\Big{\}},$ (6)
$\mathcal{X}$ $=[\bm{x}_{1}^{b},\bm{x}_{2}^{b},...,\bm{x}_{t}^{b},\bm{\lambda}_{1},\bm{\lambda}_{2},\cdots,\bm{\lambda}_{N}],$
where $\bm{x}_{t}^{b}$ denotes the body poses in the sliding window and $\bm{\lambda}_{c}$ denotes inverse-depth parameterized landmarks observed in camera $c$. $\mathcal{F}_{i}$ is the set of reliable features in $i$-th camera, $\bm{r}_{0}$ is the marginalization factor, $\bm{r}_{v_{(i,j)}}$ is the $j$-th landmark reprojection residual in the $i$-th camera. $\bm{\rho}(\cdot)$ implements Huber robustification. The Ceres solver’s Dogleg algorithm [31] efficiently handles this nonlinear least-squares problem through adaptive trust region management and analytical Jacobians derived from [1].
Scale Correction. While the iterative bundle adjustment process continuously refines body poses and feature depths through nonlinear optimization, inherent scale ambiguity leads to progressive scale drift during map expansion. To mitigate this critical issue, we implement a hierarchical scale correction module that operates during the backend optimization. Subsequently, we compute the optimal scale factors $s$ through the constrained optimization framework detailed in Section III-C. Rather than applying direct scale transformation to pose trajectories, we optimally adjust inverse feature depths based on scale factors within each camera’s coordinate system. This differential depth propagation approach enables implicit scale rectification through subsequent state estimation updates while maintaining the temporal consistency of the SLAM system.